OpenStudy (anonymous):

Factor the polynomial: 12x^2 - 5x - 3

OpenStudy (anonymous):

12x^2 -9x +4x -3 = 3x(4x-3) +1(4x-3) =(3x+1)(4x-3) =

OpenStudy (anonymous):

ty